# Designing Beyond Language: Sociotechnical Barriers in AI Health Technologies for Limited English Proficiency

### v1

@@ 3.2.1. Interest survey
Interested patient navigators completed an online screening survey that asked for demographic information, including age range, gender, ethnicity, occupation, organization, years of relevant experience in their roles, and languages spoken. We also adapted a 5-question basic AI literacy questionnaire (Grassini, 2024) in the interest form (see Section A.1). After completing the survey questions, participants signed up for a 60-minute interview within the next week. A copy of the consent form was included for the participants to review and sign before the interview.
Table 1. Storyboard themes with motivating questions and supporting literature.:
@@
Language & Communication Can AI support accurate and productive conversations? Can AI capture cultural and semantic nuances? (Alhanai et al., 2025), (Antoniak et al., 2024), (Genovese et al., 2024), (Guo et al., 2021), (Kambhamettu et al., 2025), (Liu et al., 2025b), (Mehandru et al., 2023), (Zaretsky et al., 2024)
Relationality & Comfort Can patients interact with AI comfortably? Can AI cater to the patient’s specific needs and background? (Basoah et al., 2025), (Hofmann et al., 2024), (Liu et al., 2025b), (Shaikh et al., 2024), (Wu et al., 2024)
@@
3.2.2. Storyboard interviews
Each interview began with asking about the navigator’s background, job responsibilities, and the characteristics of the patients they typically supported. We also asked about challenges they face in their role and challenges their patients face in order to identify pain points in current practices and relationships between Spanish-speaking LEP patients, navigators, and providers as opportunities for AI intervention. We then extended these discussions to consider possible outcomes of integrating AI tools in patient care.
Participants exhibited a range of knowledge on AI and ways in which they interacted with LEP individuals. To establish common ground across participants (Davidoff et al., 2007), we structured our interviews around six pre-defined storyboards. Storyboards can solicit immediate reactions and underlying needs probed by the hypothetical scenarios depicted (Holstein et al., 2019; Davidoff et al., 2007; Kawakami et al., 2023). The storyboards surfaced participants’ reactions and concerns towards potential AI systems, use cases, and patient responses that they might not otherwise think of. We explored the use of AI systems in various stages of care (e.g. before, during, and after provider visits) and in various capacities (e.g. translating, explaining, comforting). For each storyboard, we aimed to uncover in-depth understandings behind participants’ initial reactions.
@@ 3.3. Storyboard Construction
Following prior health-focused studies in HCI (Zhang and Lee, 2025; Kawakami et al., 2023), we constructed the storyboards using an iterative method drawing from existing literature and our research team’s expertise (Section 3.7). We collected and read 34 papers across HCI, AI, and medical journals that propose use cases, risks, and benefits of AI in healthcare and for LEP individuals in particular. A clinical psychologist and expert in health equity for under-served populations, in particular Spanish-speaking communities, reviewed proposed storyboards to ensure they were comprehensive, realistic, and would elicit meaningful discussions. We consolidated storyboards in six dimensions covering possible applications of AI health technologies for LEP patients (see Table 1). These dimensions included AI performance (linguistic nuances and reliability of information), ease of use (for patients and within clinic workflows), and systemic factors (access to technology and ethical concerns). Each dimension was used to create a 4-panel storyboard scenario that followed the same general structure: 1) patient encounters a challenge, 2) AI tool is introduced as a potential solution, 3) effects (either positive or negative) of AI emerge, and 4) patient reflects on the AI interaction. We strove to depict realistic AI based current commercial or prototype systems (e.g. translation systems, chatbots). We intended for the scenarios to have a mix of positive and negative outcomes to encourage participants to consider the tensions between the benefits and risks of AI for LEP patient care. Two example storyboards are shown in Figure 1, and all storyboards can be found in Section A.2. A version of each storyboard with English translations was available for participants with limited Spanish proficiency.
(a) Language & Communication theme: AI mistranslations of idioms and medical jargon showcase potential shortcomings.:
(b) Clinical Integration & Support theme: AI translator can support patient care when clinics are understaffed.:
Figure 1. Example storyboards used in our interviews.:
@@ 3.5. Participants
We recruited 14 patient navigators (8 female, 6 male) who supported Spanish-speaking LEP patients in the U.S. Participant roles included patient advocates, health liaisons, caseworkers, medical students, and interpreters. Participants came from eight distinct organizations, including medical schools, local community non-profits, and public school districts. These organizations represented diverse medical care practices and resources. For example, some organizations allowed the use of ad hoc interpreters, while others required strictly professional interpreters. Some clinics had access to interpreters for low-resource languages like Q’anjob’al, while other clinics struggled to access even common languages like Spanish. Although we did not intentionally recruit patient navigators with similar cultural backgrounds as Spanish-speaking LEP individuals, 12 out of 14 participants reported speaking English and Spanish as their primary languages and having strong ties to the communities as first- or second-generation immigrants. See Table 2 for a full breakdown of participants’ demographics and roles.
@@
Though not a requirement to participate in the study, participants also exhibited fairly high degrees of familiarity and proficiency with AI technologies. Participants reported understanding basic AI concepts $(M=3.86,SD=0.86)$ and being able to judge the benefits and risks of AI $(M=3.64,SD=0.74)$. All participants had extensive experience using AI tools such as ChatGPT or Google Translate in their work.
@@
Table 2. Summary of participants, including Occupation, Age, Sex, Race, Languages Spoken, and Years Working with Spanish-Speaking Patients (Yrs. w/ SSP).:
ID Occupation Age Sex Ethnicity/Race Languages Spoken Yrs. w/ SSP
@@
P1 Health Program Coordinator 25-34 Female Hispanic/Latinx English, Spanish 1-3 years
@@
P2 Patient Advocate 18-24 Female Hispanic/Latinx English, Spanish 3-5 years
@@
P3 Caseworker 35-44 Male Black/African American English, Spanish, French ¡1 year
@@
P4 Community Worker 18-24 Female Black/African American English, French, Lingala, Swahili ¡1 year
@@
P5 Community Navigator Volunteer 25-34 Female Hispanic/Latinx English, Spanish 5-10 years
P6 Home Visitor 45-54 Female White English, Spanish 15+ years
P7 Parent Liaison 25-34 Male Hispanic/Latinx English, Spanish ¡1 years
@@
P8 Community Health Worker 35-44 Male Hispanic/Latinx English, Spanish 15+ years
P9 Medical Student 35-44 Male Hispanic/Latinx English, Spanish 3-5 years
@@
P10 Latino Family Liaison 55-64 Female Hispanic/Latinx English, Spanish 15+ years
@@
P11 Medical Student 18-24 Male Hispanic/Latinx & White English, Spanish 3-5 years
@@
P12 Digital Literacy Program Coordinator 18-24 Female Hispanic/Latinx English, Spanish 1-3 years
@@
P13 Social Worker 45-54 Female Hispanic/Latinx English, Spanish 15+ years
@@
P14 Digital Literacy Program Coordinator 18-24 Male Hispanic/Latinx English, Spanish 1-3 years
@@ 3.6. Data Collection & Analysis
All interviews were conducted, recorded, and transcribed via Zoom. One author reviewed each transcript for accuracy, removed pauses and filler words, and paraphrased quotations cited in this paper. We followed Braun and Clark’s reflexive thematic analysis approach (Braun and Clarke, 2006, 2019). To familiarize ourselves with the data, two authors coded a subset of the interview transcripts independently. Each researcher identified quotes that were associated with our research questions and grouped these quotes with short descriptions. The two authors met to share initial ideas of interest and quotes, prompting a discussion of emerging trends that developed further into preliminary themes. These discussion included the rest of the research team in weekly and ad-hoc group meetings over the course of two weeks. The first author then reviewed the themes against the coded transcripts to ensure they were representative of the original codes. Following this consistency check, the first author coded the remaining interview transcripts.
Given the sensitive nature of the study, we anonymized participant responses, and removed personally identifiable information during transcript review and analysis. All quotes we report have been paraphrased.
@@ 3.7. Positionality Statement
Our research team is based in the U.S. and comprises researchers holding many identities. We collectively have experience in HCI, digital health, AI ethics, and clinical psychology. We were cautious that our background in HCI and AI might lead us to a solutionist perspective in this study (i.e., that AI should be applied to overcome barriers), which we actively tried to avoid by creating storyboards rather than prototypes and focusing on potential risks of AI in addition to opportunities. The team also includes a native Spanish-speaking clinical psychologist who has experience working with LEP populations, providing insight into the framing and content of the storyboards. Her personal connection with two community organizations aided with recruitment (Section 3.4), establishing familiarity and trust in our research team for those participants. The interviewing researcher is a second-generation immigrant with professional proficiency in Spanish, helping to establish some trust and familiarity with participants. However, as this researcher does not come from a similar cultural background as the patient navigators or Spanish-speaking LEP population, this difference may have influenced what participants chose to disclose and what the researcher picked up from participant responses.
@@ 4. Results
In this section, we cover our findings that address our two research questions. We start by highlighting how LEP individuals’ access to health technologies can be impeded by linguistic (Section 4.1), cultural (Section 4.2) and digital (Section 4.3 & 4.4) divides. We then describe how patient navigators saw the possible applications (Section 4.5) and risks (Section 4.6) of AI tools for LEP individuals. In this section and the following, we use the term clients to refer to the Spanish-speaking LEP individuals that the patient navigators work with, reflecting the terminology most commonly used by the navigators themselves.
@@ 4.1. Linguistic variation complicated communication surrounding care (RQ1)
Patient navigators consistently emphasized that being non-native English speakers imposed communication challenges on clients when accessing care. However, these language barriers included linguistic variation that went beyond the support of translators (human or AI-powered).
Within native Spanish speakers, dialect and colloquial differences between various countries or regions sometimes created misunderstandings (P9, P10, P12). Furthermore, merely being from a country where Spanish is the official language did not guarantee that an individual was a native Spanish speaker. Some navigators reported frequently supporting individuals from Guatemala, where the official and most spoken language is Spanish, but where there are also over 20 Mayan languages spoken, such as Q’anjob’al or K’iche’. As a native Spanish speaker from Mexico, P10 recounts an experience while supporting her students:
Kids from Guatemala and Honduras have words I’ve never heard before. Sometimes they try to explain it, and it becomes even more confusing. One of the kids described his condition as ”falling asleep,” and I was trying to explain seizures, but they have a different word for seizures in his language from Guatemala…Even when we all speak Spanish, we all have different words.
This linguistic variation posed challenges for medical interpretation, as most Spanish interpreters that the navigators and their patients encountered were from Mexico and thus unfamiliar with other dialects or languages. P10 explained that individuals ”speaking Q’anjob’al or other Guatemalan languages often struggle, since phone interpreters use words they can’t understand.” Patients recognized the limitations of interpretation services. For example, when asked for language preferences, Guatemalan individuals often default to Spanish despite inevitable misunderstandings because Spanish interpreters were easier to access (P3, P5).
The context of conversations also played a role in the nuances and connotations of words or expressions. P9 explained that although medical terminology is generally universal, ”words like ’stomach’, ’chest’, ’shoulder’, or ’neck’ can vary slightly. Sometimes translators use words from different dialects or formalities, which can confuse patients. Small differences can sometimes be the issue for patients.” However, navigators generally agreed that patients still understand the main points even if specific words differ slightly. The dialect mismatches contributed more to friction that reduced their comfort and ease of interaction with the medical system.
@@
Navigators pointed to cultural beliefs and values of their clients that clashed with U.S. healthcare practices and western medicine in general. P10 described her students’ distrust of the U.S. health system, saying that ”even students who have been here for many years don’t go to the doctor here. They talk to their grandmas or their moms in Guatemala…and their family sends them medicine.” P10 also recalled a particular instance in which the cultural belief of brujería (witchcraft) of her student from Guatemala obstructed his opportunity to receive care in the United States:
He kept having horrible cramps in his legs at night and spine pain. We went to the doctor, who recommended x-rays and additional tests, but he never wanted to go. He said he knew what he had: a problem caused by a type of demon in Guatemala, and the only cure was to see a curandero there. This is more common than I would like to see. Many kids say, ”I went to the doctor, and they say I don’t have anything, but I know I have it, and they cannot see it because they’re not from my country. Only people from my country can see it.”
Many clients preferred to follow the medical traditions from their native cultures, which may sometimes be at odds with healthcare practices recommended in the U.S. Clients often refrained from disclosing their home practices to their providers of differing backgrounds due to fear of being judged (P5), thus increasing the disconnect between providers and their knowledge of cultural practices. P6 highlights the following example around sleep practices for babies:
Doctors recommend cribs, but families often co-sleep. They’ll usually tell the doctor they use cribs because they know it’s the expected answer. If I’m there, I might say, “If they’re co-sleeping, what should they keep in mind?” That way the doctor can offer safer practices.

### v2

@@ 3.2.1. Interest survey
Interested patient navigators completed an online screening survey that asked for demographic information, including age range, gender, ethnicity, occupation, organization, years of relevant experience in their roles, and languages spoken. We also adapted a 5-question basic AI literacy questionnaire (Grassini, 2024) in the interest form (see Appendix A). After completing the survey questions, participants signed up for a 60-minute interview within the next week. A copy of the consent form was included for the participants to review and sign before the interview.
3.2.2. Storyboard interviews
Each interview began with questions about the navigator’s background, job responsibilities, and the characteristics of the patients they typically supported. We also asked about challenges they face in their role and challenges their patients face in order to identify pain points in current practices and relationships between Spanish-speaking LEP patients, navigators, and providers as opportunities for AI intervention. We then extended these discussions to consider possible outcomes of integrating AI tools in patient care.
Participants exhibited a range of knowledge on AI and ways in which they interacted with LEP individuals. To establish common ground across participants (Davidoff et al., 2007), we structured our interviews around six pre-defined storyboards. Storyboards can solicit immediate reactions and underlying needs probed by the hypothetical scenarios depicted (Holstein et al., 2019; Davidoff et al., 2007; Kawakami et al., 2023). These storyboards surfaced participants’ reactions and concerns towards potential AI systems, use cases, and patient responses that they might not otherwise think of. We explored the use of AI systems in various stages of care (e.g. before, during, and after provider visits) and in various capacities (e.g. translating, explaining, comforting).
Table 1. Storyboard themes with motivating questions and supporting literature.
A table showing 6 rows for the storyboard themes around AI health technology development, and 3 columns for dimension, motivating ideas, and relevant literature
@@ 3.2.2. Storyboard interviews
Language & Communication Can AI support accurate and productive conversations? Can AI capture cultural and semantic nuances? (Alhanai et al., 2025), (Antoniak et al., 2024), (Genovese et al., 2024), (Guo et al., 2021), (Kambhamettu et al., 2025), (Liu et al., 2025), (Mehandru et al., 2023), (Zaretsky et al., 2024)
mygrayRelationality & Comfort Can patients interact with AI comfortably? Can AI cater to the patient’s specific needs and background? (Basoah et al., 2025), (Hofmann et al., 2024), (Liu et al., 2025), (Shaikh et al., 2024), (Wu et al., 2024)
@@
black
@@ 3.3. Storyboard Construction
Following prior health-focused studies in HCI (Zhang and Lee, 2025; Kawakami et al., 2023), we constructed the storyboards using an iterative method drawing from existing literature and our research team’s expertise (Section 3.7). We synthesized 34 papers across HCI, AI, and medical journals that propose use cases, risks, and benefits of AI in healthcare and for LEP individuals in particular. Then, a clinical psychologist and expert in health equity for under-served populations, in particular Spanish-speaking communities, reviewed proposed storyboards to ensure they were comprehensive, realistic, and would elicit meaningful discussions. We consolidated storyboards in six dimensions covering possible applications of AI health technologies for LEP patients (see Section 3.2.2). These dimensions included AI performance (linguistic nuances and reliability of information), ease of use (for patients and within clinical workflows), and systemic factors (access to technology and ethical concerns). Each dimension was used to create a 4-panel storyboard scenario that followed the same general structure: 1) patient encounters a challenge, 2) AI tool is introduced as a potential solution or support, 3) effects (either positive or negative) of AI emerge, and 4) patient reflects on the interaction with AI. Two example storyboards are shown in Figure 1, and all storyboards are provided in the supplementary materials. A version of each storyboard with English translations was available for participants with limited Spanish proficiency.
All storyboards depict patients as the direct users of AI (except for the ”Language & Communication” theme in which a patient and navigator interact with AI together) because we initially centered this work around AI design considerations for LEP patients in isolation since this group is often left out of technology design. However, navigators later revealed barriers that complicate this reality, as further discussed in Section 5.1 and Section 5.3. Furthermore, we intended for the scenarios to have a mix of positive and negative outcomes to encourage participants to consider the tensions between the benefits and risks of AI for LEP patient care. We strove to depict realistic AI-based current commercial or prototype systems (e.g. translation systems, chatbots). However, since storyboards inherently depict imagined rather than observed behavior, we intended to use them to elicit expectations, concerns, and situated judgments. This limitation is further discussed in Section 5.4.
(a) Language & Communication theme: AI mistranslations of idioms and medical jargon showcase potential shortcomings.
A storyboard with four panels that show: 1) A woman is experiencing chest pains, 2) A chatbot interface on a computer screen shows a conversation in which the woman asks the chatbot in Spanish how to explain her symptoms in English, and the chatbot provides a translation, 3) A patient advocate notices and corrects mistranslations, 4) The woman and her patient advocate practice rehearsing her symptoms in English.
(b) Clinical Integration & Support theme: AI translator can support patient care when clinics are understaffed.
A storyboard with four panels that show: 1) A crowded clinic waiting room, 2) A Spanish-speaking patient is conversing with an English-speaking provider using an AI app to translate, 3) The patient and provider continue conversing through the AI translator, 4) The patient leaves the appointment and continues to use the AI app for more information.
Figure 1. Example storyboards used in our interviews.
Two 4-panel digitally-illustrated cartoon storyboards. The dialogue in the storyboards is in Spanish, but each panel has an associated English caption.
@@ 3.5. Participants
We recruited 14 patient navigators (8 female, 6 male) who supported Spanish-speaking LEP patients in the U.S. Participant roles included patient advocates, health liaisons, caseworkers, medical students, and interpreters. Participants came from eight distinct organizations, including medical schools, local community non-profits, and public school districts. Most of these organizations specifically serve, or at least have strong ties to, the Spanish-speaking or broader immigrant communities in their local areas. Therefore, patients either actively seek support from these organizations or are referred to them by word-of-mouth. The communities that these 14 participants work with generally consist of first- or second-generation immigrants from Hispanic countries. Because many of these LEP immigrants are undocumented, tightened immigration policies in the U.S. have led to greater institutional distrust and fear that disclosing their personal information may lead back to immigration enforcements. Therefore, navigators spent considerable efforts helping these individuals overcome barriers due to their status, such as applying for health insurance without a social security number.
@@
These navigators also represent organizations with diverse medical care practices and resources. For example, some organizations allowed the use of ad hoc interpreters, while others required strictly professional interpreters. Some clinics had access to interpreters for low-resource languages like Q’anjob’al, while other clinics struggled to access even common languages like Spanish. Although we did not intentionally recruit patient navigators with similar cultural backgrounds as Spanish-speaking LEP individuals, 12 out of 14 participants reported speaking English and Spanish as their primary languages and having strong ties to the communities as first- or second-generation immigrants. We conducted interviews in English to align with participants’ professional working language and to reduce risks of misinterpretation in analysis. See Table 2 for a full breakdown of participants’ demographics and roles.
@@
Though not a requirement to participate in the study, participants also exhibited fairly high degrees of familiarity and proficiency with AI technologies. On a 5-point Likert scale ranging from 1 (strongly disagree) to 5 (strongly agree), participants reported understanding basic AI concepts $(M=3.86,SD=0.86)$ and being able to judge the benefits and risks of AI $(M=3.64,SD=0.74)$. All participants had extensive experience using AI tools such as ChatGPT or Google Translate in their work.
@@
Table 2. Summary of participants, including Occupation, Age, Gender, Race, Languages Spoken, and Years Working with Spanish-Speaking Patients (Yrs. w/ SSP).
A table showing demographic information of 14 participants, with columns for participant, occupation, age range, gender, ethnicity/race, languages spoken, and years working with Spanish-speaking patients.
ID Occupation Age Gender Ethnicity/Race Languages Spoken Yrs. w/ SSP
@@
P1 Health Program Coordinator 25-34 Woman Hispanic/Latinx English, Spanish 1-3 years
@@
P2 Patient Advocate 18-24 Woman Hispanic/Latinx English, Spanish 3-5 years
@@
P3 Caseworker 35-44 Man Black/African American English, Spanish, French ¡1 year
@@
P4 Community Worker 18-24 Woman Black/African American English, French, Lingala, Swahili ¡1 year
@@
P5 Community Navigator Volunteer 25-34 Woman Hispanic/Latinx English, Spanish 5-10 years
P6 Home Visitor 45-54 Woman White English, Spanish 15+ years
P7 Parent Liaison 25-34 Man Hispanic/Latinx English, Spanish ¡1 years
@@
P8 Community Health Worker 35-44 Man Hispanic/Latinx English, Spanish 15+ years
P9 Medical Student 35-44 Man Hispanic/Latinx English, Spanish 3-5 years
@@
P10 Latino Family Liaison 55-64 Woman Hispanic/Latinx English, Spanish 15+ years
@@
P11 Medical Student 18-24 Man Hispanic/Latinx & White English, Spanish 3-5 years
@@
P12 Digital Literacy Program Coordinator 18-24 Woman Hispanic/Latinx English, Spanish 1-3 years
@@
P13 Social Worker 45-54 Woman Hispanic/Latinx English, Spanish 15+ years
@@
P14 Digital Literacy Program Coordinator 18-24 Man Hispanic/Latinx English, Spanish 1-3 years
@@ 3.6. Data Collection & Analysis
All interviews were conducted, recorded, and transcribed via Zoom. One author reviewed each transcript for accuracy, removed pauses and filler words, and paraphrased quotations cited in this paper. We followed Braun and Clark’s reflexive thematic analysis approach (Braun and Clarke, 2006, 2019). To familiarize ourselves with the data, two authors coded a subset of the interview transcripts independently. Each researcher identified quotes that were associated with our research questions and grouped these quotes with short descriptions. The two authors met to share initial ideas of interest and quotes, prompting a discussion of emerging trends that developed further into preliminary themes. These discussions included the rest of the research team in weekly and ad-hoc group meetings over the course of two weeks. The first author then reviewed the themes against the coded transcripts to ensure they were representative of the original codes. Following this consistency check, the first author coded the remaining interview transcripts.
Given the sensitive nature of the study, we anonymized participant responses, and removed all personally identifiable information during transcript review and analysis. All quotes we report have been paraphrased to remove any personally identifiable information while preserving the semantic content and intent of participants’ statements. We retained key linguistic markers (e.g., references to dialect mismatch, idioms) when essential to the analytic claim.
@@ 3.7. Positionality Statement
Our research team is based in the U.S. and comprises researchers of many identities. We collectively have experience in HCI, digital health, AI ethics, and clinical psychology. We were cautious that our background in HCI and AI might lead us to a solutionist perspective in this study (i.e., that AI should be applied to overcome barriers), which we actively tried to avoid by creating storyboards rather than prototypes and focusing on potential risks of AI in addition to opportunities. The team also includes an immigrant, native Spanish-speaking clinical psychologist who has lived experience, expertise, and clinical experience working with LEP populations, providing insight into the framing and content of the storyboards. Her personal connection with two community organizations aided with recruitment (Section 3.4), establishing familiarity and trust in our research team for those participants. The interviewing researcher is a second-generation immigrant with professional proficiency in Spanish, helping to establish some trust and familiarity with the participants. However, as this researcher does not come from a similar cultural background as the patient navigators or Spanish-speaking LEP population, this difference may have influenced what participants chose to disclose and what the researcher picked up from participant responses.
@@ 4. Results
In this section, we cover our findings that address our two research questions. We start by highlighting how LEP individuals’ access to health technologies can be impeded by linguistic (Section 4.1), cultural (Section 4.2), and digital (Section 4.3 & 4.4) divides. We then describe how patient navigators saw the possible applications (Section 4.5) and risks (Section 4.6) of AI tools for LEP individuals. In this section and the following, we use the term clients to refer to the Spanish-speaking LEP individuals that the patient navigators work with, reflecting the terminology most commonly used by the navigators themselves. Section 4 summarizes our findings of the barriers that LEP individuals face (RQ1, Section 4.1-Section 4.4), and Section 4.4 summarizes our findings of the risks and opportunities for AI to support health communication (RQ2, Section 4.5-Section 4.6). Both tables also include associated AI design considerations (Section 5).
Table 3. Summary of barriers LEP individuals face with respect to health technologies (RQ1), how those barriers manifest, and associated design implications for AI. Bullet points map observed challenges to a potential implication for AI design.
Barrier Observed Challenges Implications for AI
mygray      Cultural differences Traditional or home remedies may conflict with western medical advice Fear of being judged prevents full disclosure Respect for authority discourages contradicting or asking questions Acknowledge cultural practices and values while offering practical steps for care Capture nonverbal cues that signify discomfort Avoid didactic or dismissive tones that misalign advice by ignoring cultural nuances
Literacy Low reading literacy in English and/or Spanish Low digital literacy due to limited experience with usable technology (phones, computers, Internet) Lack of specialized medical knowledge complicates understanding health information Offer voice- or picture-based controls Accompany AI with resources to teach prompting and information validation skills; build simpler models that do not require state-of-the-art devices Provide simplified explanations that do not require extensive prior medical knowledge
Privacy Reluctance to provide personal information Fear that personal data could end up in unwanted entities Hesitancy to engage with digital systems, preferring in-person means Access AI services without requiring personal information; anonymize when possible Minimize storing data and history; offer options to anonymize Build models that run locally on-device, without requiring external data transmission
black
A table showing 4 rows of barriers that patients face with respect to health technologies. Columns include barriers, observed challenges, and implications for AI.
@@ 4.1. Linguistic variation complicated communication surrounding care (RQ1)
All patient navigators consistently emphasized that being non-native English speakers imposed communication challenges on clients when accessing care. However, these language barriers included linguistic variation that went beyond the support of translators (human or AI-powered).
Within native Spanish speakers, dialectal and colloquial differences between various countries or regions sometimes created misunderstandings (P9, P10, P12). Furthermore, merely being from a country where Spanish is the official language did not guarantee that an individual spoke Spanish. Some navigators reported supporting individuals from Guatemala, where the official and most spoken language is Spanish, but where there are also over 20 Mayan languages spoken, such as Q’anjob’al or K’iche’. As a native Spanish speaker from Mexico, P10 recounts an experience while supporting her students:
”Kids from Guatemala and Honduras have words I’ve never heard before. Sometimes they try to explain it, and it becomes even more confusing. One of the kids described his condition as ‘falling asleep,’ and I was trying to explain seizures, but they have a different word for seizures in his language from Guatemala…Even when we all speak Spanish, we all have different words.”
This linguistic variation posed challenges for medical interpretation, as most Spanish interpreters that the navigators and their patients encountered were from Mexico and thus less familiar with other dialects or languages. P10 explained that individuals ”speaking Q’anjob’al or other Guatemalan languages often struggle, since phone interpreters use words they can’t understand.” Patients recognized the limitations of interpretation services. For example, when asked for language preferences, Guatemalan individuals often default to Spanish despite inevitable misunderstandings because Spanish interpreters were easier to access (P3, P5).
The context of conversations also played a role in the nuances and connotations of words or expressions. P9 explained that although medical terminology is generally universal, ”words like ‘stomach’, ‘chest’, ‘shoulder’, or ‘neck’ can vary slightly. Sometimes translators use words from different dialects or formalities, which can confuse patients. Small differences can sometimes be the issue for patients.” However, navigators generally agreed that patients still understand the main points even if specific words differ slightly. The dialect mismatches contributed more to friction that reduced their comfort and ease of interaction with the medical system.
@@
Navigators pointed to clients’ cultural beliefs and values that clashed with U.S. healthcare practices and western medicine. P10 described her students’ distrust of the U.S. health system, saying that ”even students who have been here for many years don’t go to the doctor here. They talk to their grandmas or their moms in Guatemala…and their family sends them medicine.” P10 also recalled a particular instance in which the cultural belief of brujería (witchcraft) of her student from Guatemala obstructed his opportunity to receive care in the United States:
”He kept having horrible cramps in his legs at night and spine pain. We went to the doctor, who recommended x-rays and additional tests, but he never wanted to go. He said he knew what he had: a problem caused by a type of demon in Guatemala, and the only cure was to see a curandero there. This is more common than I would like to see. Many kids say, ‘I went to the doctor, and they say I don’t have anything, but I know I have it, and they cannot see it because they’re not from my country. Only people from my country can see it.’”
Many clients preferred to follow the medical traditions from their native cultures, which may sometimes be at odds with healthcare practices recommended in the U.S. Clients often refrained from disclosing their home practices to providers of differing backgrounds due to fear of being judged (P5), thus increasing the disconnect between providers and their knowledge of cultural practices. P6 highlighted an example around sleep practices for babies:
”Doctors recommend cribs, but families often co-sleep. They’ll usually tell the doctor they use cribs because they know it’s the expected answer. If I’m there, I might say, ‘If they’re co-sleeping, what should they keep in mind?’ That way the doctor can offer safer practices.”